OpenStudy (ashlynxo):

Use the discriminant to find the number of real solutions of the equation 3x^2 – 5x + 4 = 0. Show your work.

OpenStudy (anonymous):

I forgot what is discriminant of a quadratic. Do you remember?

OpenStudy (ashlynxo):

Well, the textbook definition is "the discriminant is the expression under the radical sign in the quadratic formula."

OpenStudy (johnweldon1993):

From the quadratic formula.. \[\large \frac{-b \pm \color\red{\sqrt{b^2 - 4ac}}}{2a}\] The red is your descriminant

OpenStudy (johnweldon1993):

So from your equation \[\large 3x^2-5x + 4 = 0\] A,B and C are the coefficients of the quadratic 3 = A -5 = B 4 = C so we plug those into the discriminant and what do we get?